OpenStudy (anonymous):

A car completes a journey in 10 minutes. For the first half of the distance the speed was 60 km/h and for the second half the speed was 40 km/h. How far is the journey?

OpenStudy (anonymous):

8 kms

OpenStudy (zzr0ck3r):

\(10\) minutes is \(\frac{1}{6}\) of an hour These next two lines follow from distance=rate\(\times\)time and thus time = distance over rate i.e. \(d=r*t\implies t=\frac{d}{r}\) The time traveled at \(60\) is \(\frac{1}{2}\) the distance over \(60\) kmh The time traveled at \(50\) is \(\frac{1}{2}\) the distance over \(40\) kmh Now we know the total time is \(\frac{1}{6}\) of an hour, so we should add our first two times from above and that should be equal to \(\frac{1}{6}\) \(\frac{1}{6}=\frac{1}{2}*d*\frac{1}{60}+\frac{1}{2}*d*\frac{1}{40}\) Solve for distance \(d\).

OpenStudy (anonymous):

Thank you zr0ck!